# Qwen2.5-Coder Technical Report

### v1

@@ 7.2 Code Reasoning
To assess the code reasoning capabilities of the Qwen2.5-Coder series instruct models, we performed an evaluation on CRUXEval (Gu et al., 2024). As illustrated by the experimental results in Table 15, the Qwen2.5-Coder-7B-Instruct model achieved Input-CoT and Output-CoT accuracies of 65.8% and 65.9%, respectively. This represents a notable improvement over the DS-Coder-V2-Lite-Instruct model, with gains of 12.8% in Input-CoT accuracy and 13.0% in Output-CoT accuracy. Furthermore, the Qwen2.5-Coder-7B-Instruct model outperformed larger models, such as the CodeStral-22B and DS-Coder-33B-Instruct, underscoring its superior code reasoning capabilities despite its smaller size.
Figure 7 illustrates the relationship between model sizes and code reasoning capabilities. The Qwen2.5-Coder instruct models stand out for delivering superior code reasoning performance with the fewest parameters, surpassing the results of other open-source large language models by a significant margin. According to this trend, we expect that code reasoning performance comparable to GPT-4o could be achieved with a model around the 30 billion parameters scale.
@@
DS-Coder-1.3B-Instruct 1.3B 17.3* 21.1*
Qwen2.5-Coder-1.5B-Instruct 1.5B 30.1 33.1
@@
DS-Coder-6.7B-Instruct 6.7B 34.6* 42.1*
DS-Coder-V2-Lite-Instruct 2.4/16B 48.9* 55.6*
CodeQwen1.5-7B-Chat 7B 23.3 33.1
Qwen2.5-Coder-7B-Instruct 7B 50.4 57.1
@@
CodeStral-22B 22B 35.3* 45.1*
DS-Coder-33B-Instruct 33B 49.6* 59.4*
DS-Coder-V2-Instruct 21/236B 73.7 -
Table 16: The code editing ability of different instruct models evaluated by Aider benchmark. * indicates that the experimental results have been reproduced in our experiments, and the whole edit-format was consistently applied across all experiments.
@@ 7.3 Code Editing
Aider [^8] https://github.com/paul-gauthier/aider has created a code editing benchmark designed to quantitatively measure its collaboration with large language models (LLMs). Drawing from a set of 133 Python exercises sourced from Exercism [^9] https://github.com/exercism/python, the benchmark tests the ability of Aider and LLMs to interpret natural language programming requests and translate them into executable code that successfully passes unit tests. This assessment goes beyond evaluating raw coding proficiency; it also examines how effectively LLMs can edit existing code and format those modifications for seamless integration with Aider’s system, ensuring that local source files can be updated without issues. The comprehensive nature of this benchmark reflects both the technical aptitude of the LLMs and their consistency in task completion.
Table 16 highlights the performance of several language models in the Code Editing task. Among them, Qwen2.5-Coder-7B-Instruct demonstrates outstanding code repair capabilities. Despite its relatively modest size of 7 billion parameters, it achieves an impressive PASS@1 accuracy of 50.4%, significantly outperforming comparable models. Notably, it also surpasses larger models such as CodeStral-22B (22 billion parameters) and DS-Coder-33B-Instruct (33 billion parameters), showcasing its remarkable efficiency and effectiveness in code editing tasks.
@@ 7.4 Text-to-SQL
Prompt template for text-to-SQL tasks.
Figure 8: Prompt template of Qwen2.5-Coder for text-to-SQL tasks.
Thanks to the use of finely crafted synthetic data during both pre-training and fine-tuning, we significantly enhanced Qwen2.5-Coder’s capability in Text-to-SQL tasks. We selected two well-known benchmarks, Spider (Yu et al., 2018) and BIRD (Li et al., 2024), for comprehensive evaluation. To ensure a fair comparison between Qwen2.5-Coder and other open-source language models on this task, we used a unified prompt template as input, following the work of Chang & Fosler-Lussier (2023). As shown in Figure 8, the prompt consists of table representations aligned with database instructions, examples of table content, optional additional knowledge, and natural language questions. This standardized prompt template minimizes biases that may arise from prompt variations. As shown in Figure 9, Qwen2.5-Coder outperforms other code models of the same size on the Text-to-SQL task.
@@ 7.5 Math Reasoning and General Natural Language
In this section, we present a comparative analysis of the performance between our Qwen2.5-Coder-7B-Instruct model and the DS-Coder-V2-Lite-Instruct model, focusing on both mathematical computation and general natural language processing tasks. As indicated in Table 17, the Qwen2.5-Coder-7B-Instruct model outperforms the DS-Coder-V2-Lite-Instruct in 11 out of 12 tasks. This result underscores the model’s versatility, excelling not only in complex coding tasks but also in sophisticated general tasks, thus distinguishing it from its competitors.
@@ 8 Conclusion
This work introduces Qwen2.5-Coder, the latest addition to the Qwen series. Built upon Qwen2.5, a top-tier open-source LLM, Qwen2.5-Coder has been developed through extensive pre-training and post-training of Qwen2.5-1.5B and Qwen2.5-7B on large-scale datasets. To ensure the quality of the pre-training data, we have curated a dataset by collecting public code data and extracting high-quality code-related content from web texts, while filtering out low-quality data using advanced classifiers. Additionally, we have constructed a meticulously designed instruction-tuning dataset to transform the base code LLM into a strong coding assistant.

### v2

@@ 7.2 Code Reasoning
To evaluate the code reasoning capabilities of the Qwen2.5-Coder series instruct models, we conducted an assessment on the CRUXEval (Gu et al., 2024) dataset. As shown in Table 18, the Qwen2.5-Coder-7B-Instruct model achieved Input-CoT and Output-CoT accuracies of 65.8% and 65.9%, respectively—demonstrating a substantial improvement over the DS-Coder-V2-Lite-Instruct model, with gains of 12.8% in Input-CoT accuracy and 13.0% in Output-CoT accuracy. Additionally, the Qwen2.5-Coder-7B-Instruct model outperformed larger models, including CodeStral-22B and DS-Coder-33B-Instruct, highlighting its advanced code reasoning capabilities despite its smaller size. Notably, our Qwen2.5-Coder-32B-Instruct model achieved accuracies of 75.2% and 83.4% on Input-CoT and Output-CoT, respectively, significantly outperforming other open-source code models (including DS-Coder-V2-Instruct) and underscoring its robust performance in code reasoning.
Figure 10 illustrates the relationship between model sizes and code reasoning capabilities. The Qwen2.5-Coder instruct models stand out for delivering superior code reasoning performance with the fewest parameters, surpassing the results of other open-source large language models by a significant margin.
@@
DS-Coder-1.3B-Instruct 1.3B 18.0 18.8
Yi-Coder-1.5B-Chat 1.5B 17.3 17.3
Qwen2.5-Coder-1.5B-Instruct 1.5B 28.6 31.6
3B+ Models
Qwen2.5-Coder-3B-Instruct 3B 33.8 39.1
@@
CodeLlama-7B-Instruct 7B 1.5 1.5
DS-Coder-6.7B-Instruct 6.7B 37.6 44.4
CodeQwen1.5-7B-Chat 7B 24.8 38.3
Yi-Coder-9B-Chat 9B 45.9 54.1
DS-Coder-V2-Lite-Instruct 2.4/16B 44.4 52.6
Qwen2.5-Coder-7B-Instruct 7B 51.9 57.9
13B+ Models
CodeLlama-13B-Instruct 13B 1.5 1.5
Starcoder2-15B-Instruct-v0.1 15B - -
Qwen2.5-Coder-14B-Instruct 14B 58.6 69.2
@@
CodeLlama-34B-Instruct 34B 12.8 15.0
CodeStral-22B-v0.1 22B 36.8 51.1
DS-Coder-33B-Instruct 33B 50.4 59.4
CodeLlama-70B-Instruct 70B 1.5 1.5
DS-Coder-V2-Instruct 21/236B - 72.9
Qwen2.5-Coder-32B-Instruct 32B 60.9 73.7
Closed-APIs
Claude-3.5-Sonnet-20240620 - 59.4 66.2
Claude-3.5-Sonnet-$2$ 71.4 86.5
GPT-4o-mini-2024-07-18 - 43.6 55.6
GPT-4o-2024-08-06 - 56.8 71.4
o1-mini - 49.6 70.7
o1-preview - 69.9 88.0
Table 19: The code editing ability of different instruct models evaluated by Aider benchmark. The whole edit-format was consistently applied across all our experiments.
Figure 11: The evaluation results on CodeEditBench.
@@ 7.3 Code Editing
Aider
Aider [^9] https://github.com/paul-gauthier/aider has created a code editing benchmark designed to quantitatively measure its collaboration with large language models (LLMs). Drawing from a set of 133 Python exercises sourced from Exercism [^10] https://github.com/exercism/python, the benchmark tests the ability of Aider and LLMs to interpret natural language programming requests and translate them into executable code that successfully passes unit tests. This assessment goes beyond evaluating raw coding proficiency; it also examines how effectively LLMs can edit existing code and format those modifications for seamless integration with Aider’s system, ensuring that local source files can be updated without issues. The comprehensive nature of this benchmark reflects both the technical aptitude of the LLMs and their consistency in task completion. Table 19 highlights the performance of several language models in the Code Editing task. Among these models, Qwen2.5-Coder-7B-Instruct exhibits exceptional code repair capabilities. Despite its relatively modest scale of 7 billion parameters, it achieves an impressive PASS@1 accuracy of 51.9%, significantly outperforming comparable models. Remarkably, it also surpasses larger models such as CodeStral-22B and DS-Coder-33B-Instruct , highlighting its remarkable efficiency and effectiveness in code editing tasks. Our Qwen2.5-Coder-32B-Instruct model achieves even higher accuracy, with Pass@1 and Pass@2 rates reaching 60.9% and 73.7%, respectively.
CodeEditorBench
An effective code assistant must excel in generating code based on given specifications, as well as in modifying or debugging existing code to meet evolving requirements or resolve issues. In evaluating Qwen2.5-Coders proficiency in code modification tasks, we focused on the CodeEditorBench (Guo et al., 2024b) suite, which assesses performance across four key dimensions: Debugging, Translation, Switching, and Polishing. We employed the same evaluation approach used in the original paper, relying on win rate as the metric for overall performance across diverse problem types. The win rate was computed for each problem category and then averaged across all categories to obtain the overall score. The results in Figure 11 show that Qwen2.5-Coder-32B-Instruct achieves a win rate comparable to DS-Coder-V2-Instruct (86.2% win rate), which features a significantly larger 236 billion parameter scale.
@@ 7.4 Text-to-SQL
Thanks to the use of finely crafted synthetic data during both pre-training and fine-tuning, we significantly enhanced Qwen2.5-Coder’s capability in Text-to-SQL tasks. We selected two well-known benchmarks, Spider (Yu et al., 2018) and BIRD (Li et al., 2024a), for comprehensive evaluation. To ensure a fair comparison between Qwen2.5-Coder and other open-source language models on this task, we used a unified prompt template as input, following the work of Chang & Fosler-Lussier (2023). The evaluation prompt consists of table representations aligned with database instructions, examples of table content, optional additional knowledge, and natural language questions. This standardized prompt template minimizes biases that may arise from prompt variations. As shown in Figure 12, Qwen2.5-Coder outperforms other code models of the same size on the Text-to-SQL task.
Figure 13: The table understanding evaluation on TableBench.
@@
Qwen2.5-Coder-14B-Instruct 14B 66.8 94.2 66.0 40.1 47.3 10.0
Qwen2.5-Coder-32B-Instruct 32B 76.4 93.0 68.3 42.5 47.7 20.0
@@
Qwen2.5-Coder-14B-Instruct 14B 50.0 71.7 55.6 66.5 66.2 36.8
Qwen2.5-Coder-32B-Instruct 32B 55.0 77.6 62.3 79.9 68.9 41.8
@@ 7.5 Math Reasoning and General Natural Language
In this section, we provide a comparative analysis of the performance between our Qwen2.5-Coder series models and the DS-Coder-V2 series models, with a focus on both mathematical computation and general natural language processing tasks. The results in Table 20 highlight the versatility of the Qwen2.5-Coder series, which excels not only in complex coding tasks but also in advanced general-purpose tasks, setting it apart from its competitors.
7.6 Table Understanding
To evaluate the understanding capabilities of structured data, we further evaluate the Qwen2.5-Coder on a comprehensive and complex benchmark TableBench (Wu et al., 2024b), which includes 18 fields within four major categories of table question answering (TableQA) capabilities. We compare Qwen2.5-Coder with other LLMs under the textual chain-of-thought (TCoT) setting. Figure 13 demonstrates that Qwen2.5-Coder-32B-Instruct gets the best performance 45.1 on TableBench.
8 Discussion: Scaling is All You Need
In Figure 14, We present a comparison of different sizes of Qwen2.5-Coder with other open-source LLMs on MBPP-3shot and LiveCodeBench. For the base LLM, we choose MBPP-3shot as the evaluation metric. Our extensive experiments show that MBPP-3shot is more suitable for evaluating base models and correlates well with the actual performance of the models. For the instruction model, we select the latest 4 months of LiveCodeBench (2024.07$\sim$2024.11) questions as the evaluation to strictly avoid test data contamination, truly reflecting the OOD capabilities of the LLM. There is a positive correlation between model size and model performance, and Qwen2.5-Coder has achieved state-of-the-art performance across all sizes, encouraging us to continue exploring larger sizes of code LLM.
Figure 14: The evaluation results of Qwen2.5-Coder models with different sizes on MBPP-3shot and LiveCodeBench.
@@ 9 Conclusion
This work introduces Qwen2.5-Coder, the latest addition to the Qwen series. Built upon Qwen2.5, a top-tier open-source LLM, Qwen2.5-Coder has been developed through extensive pre-training and post-training of Qwen2.5-0.5B/1.5B/3B/7B/14B/32B on large-scale datasets. To ensure the quality of the pre-training data, we have curated a dataset by collecting public code data and extracting high-quality code-related content from web texts, while filtering out low-quality data using advanced classifiers. Additionally, we have constructed a meticulously designed instruction-tuning dataset to transform the base code LLM into a strong coding assistant.